The coefficients of variations for the two distributions are 60 and 70 and its standard deviations are 21 and 16 respectively. Determine its arithmetic mean.

Grade:12

1 Answers

Harshit Singh
3 years ago
Dear Student

Given that,

Coefficient of Variations (C.V of 1st distribution) = 60, σ1= 21

Coefficient of Variations (C.V of 2nd distribution) = 70, σ2= 16

Let μ1and μ2are the means of the 1st and the 2nd distribution.

We know that the formula to find the arithmetic mean is given as:

Coefficient of Variations(C.V) = (Standard Deviation/arithmetic Mean) x 100

Thus, Arithmetic Mean = (Standard Deviation/C.V)x100

Therefore, the arithmetic mean for the 1st deviation is given by:

μ1= [σ1/ (c.v of 1st distribution)]x100

μ1= (21/60)x100

μ1= 0.35×100

μ1= 35

Similarly for μ2:

μ2= [σ2/ (c.v of 2nd distribution)]x100

μ2= (16/70)x100

μ2= 0.2285×100

μ2= 22.85

Thanks